A 4% solution contains 4% disinfectant and 96% water. How many ounces of solution are made with 2 ounces of disinfectant?
Answer by edjones(8007)

x/2 = 96/4
x=192/4
=48 oz water
48+2=50 oz of 4% solution.
.
Ed